A radioactive material decays to (1/16)th of its inital amount in 2hrs. What is the half life of the sample?
1
Expert's answer
2019-09-13T10:15:55-0400

N=N0*2(-t/T); N=N0/16; t=2 hours; N/N0=2(-t/T); t/T=log2(N0/N); T=t/log2(N0/N); T=0.5 hours = 30 min